1.     Find the enthalpy change for the reaction       (5 points)

         CS2 (l)  + 3 O2 (g) → CO2 (g)  + 2 SO2(g)      

When:   C(s) + O2(g) → CO2 (g)             

                       S(s) + O2 (g) → SO2(g)              

               C(s) + 2S(s) → CS2(l)       


1
Expert's answer
2022-02-06T11:20:01-0500

"\Delta{H}_{reaction}=\sum{\Delta{H_{products}-\sum\Delta{H}_{reactants}}}=-393.5+2(-296.8)-87.9=-1075kJ/mol"


Answer: -1075 kJ/mol